OpenStudy (anonymous):

Solve: x2 + 3x = 28 {-7, 4} {-4, 7} {-7, -4} {7, 4}

OpenStudy (anonymous):

change into x^2 + 3x - 28 = 0 factor out

OpenStudy (anonymous):

x^2 + 3x = 28 x^2 + 3x -28 =0 (x+7)(x-4) =0 (x+7)=0 or (x-4) =0 Can you solve it?

OpenStudy (anonymous):

Can you explain in brief for the 2nd and 3rd step

OpenStudy (anonymous):

x^2 + 3x -28 =0 x^2 + 7x - 4x - 28 =0 x(x+7) - 4(x+7) =0 (x-4)(x+7) =0 Do I make it clearer to you?

OpenStudy (anonymous):

k thanks i made a mistake sorry :/

OpenStudy (anonymous):

Welcome and no worry!